# Beyond Accuracy: A Stability-Aware Metric for Multi-Horizon Forecasting

### v1

@@ 5. Conclusion
Our metric is built upon the energy score and energy distance. It is highly customizable, as one can specify relative weights to discount forecasts at longer horizons based on problem-specific needs, and can adjust the relative importance of stability w.r.t. accuracy through the stability multiplier parameter. This makes the metric applicable across diverse forecasting contexts, from applications where stability is paramount (such as capacity planning and resource allocation) to those where accuracy at specific horizons takes precedence.
To validate the utility of the forecast AC score, we implemented it as a differentiable objective function for training seasonal ARIMA models. We performed experiments on a subset of M4 Hourly benchmark dataset including 50 time series. As a result, models trained using our metric yield substantial improvements over traditionally trained models. An average reduction of 75.4% in the standard deviation of forecasts targeting the same future timestamps is observed on the test set. This indicates that our metric is capable of training models that produce more stable forecasts.
@@ 6. Future Directions
We mention directions for extending our current work.
@@ 6.2. More Complex Forecasting Architectures
While we have demonstrated the usefulness of our metric with the seasonal ARIMA model, the metric can be applied to training more complex models as their objective functions.

### v2

@@ 5. Conclusion
Our metric is built upon the energy score and energy distance. It is highly customizable, as one can specify relative weights to discount forecasts at longer horizons based on problem-specific needs, and can adjust the relative importance of stability with respect to accuracy through the stability multiplier parameter. This makes the metric applicable across diverse forecasting contexts, from applications where stability is paramount (such as capacity planning and resource allocation) to those where accuracy at specific horizons takes precedence.
To validate the utility of the forecast AC score, we implemented it as a differentiable objective function for training SARI models. We performed experiments on the M4 Hourly benchmark dataset. As a result, models trained using our metric yield substantial improvements over traditionally trained models in terms of multi-horizon accuracy and stability. In terms of stability, the AC-optimized model generated out-of-sample forecasts with 8.9% vertical volatility relative to that of MLE-fitted model. On the accuracy side, the accuracy of forecasts targeting middle to long horizons are substantially improved for the AC-optimized model. While the one-step-ahead forecasts suffered a 7.5% degradation in MAPE, forecasts targeting all subsequent horizons experienced improved accuracy, up to 26% measured by MAPE. This indicates that our metric is capable of training models that produce more stable and accurate multi-step forecasts at relatively small cost of the one-step-ahead forecasts.
@@ 6. Future Directions
In the final section of the paper, we discuss some potential directions for extending our current work.
6.1. Application to more advanced models
Our current implementation excludes moving average (MA) terms from the SARIMA specification. This restriction addresses a computational challenge: in a differentiable implementation, MA components require forecast error history, which depends on forecasts from previous training epochs. This creates deep computational graphs that significantly increase memory requirements and training time. Future work will incorporate MA components through techniques such as error approximations to reduce computational cost.
We additionally plan to evaluate the AC score on modern forecasting models including DeepAR, Temporal Fusion Transformers, and N-BEATS. These experiments will provide a full picture of the metric’s applicability beyond classical statistical models.
@@ 6.3. More Complex Forecasting Architectures
While we have demonstrated the utility of our metric with the SARI model, the metric can also be utilized as an objective function in training more complex models.